4-3(m+1)=(-38) also if you could send other examples of this it would be greatly appreciated!
Answer by checkley75(3666)

4-3(M+1)=-38
4-3M-3=-38
-3M=-38-4+3
-3M=-39
M=-39/-3
M=13 ANSWER.
PROOF
4-3(13+1)=-38
4-3(14)=-38
4-42=-38
-38=-38
---------------------------------------------------------------
5-3(X+5)+3=8
5-3X-15+3=8
-3X=8-5+15-3
-3X=15
X=15/-3
X=-5 ANSWER.
PROOF
5-3(-5+5)+3=8
5-3*0+3=8
8=8